OpenStudy (anonymous):

Decibel question? Why is there a ratio in the Db equation?? what does this ratio represent? i thought decibels represented sound, such as a lawnmower, why does there have to be two sound/powers?? please help!

OpenStudy (anonymous):

\[\frac{ P_{2} }{ P _{1} }\] is the ratio im talking about

OpenStudy (vincent-lyon.fr):

Bels and decibels being a logarithmic scale, you need a reference power (non zero) that you call 0 on the logarithmic scale. P1 is usually taken as 10E(-12) W/m²

OpenStudy (anonymous):

I'm currently writing a project on the decibel equation and its graphic applications, but i'm corresponding it to sound, such as music (70 decibels, for example). what would be my units in the ratio? and how can i explain this ratio in my project?

OpenStudy (vincent-lyon.fr):

70dB mean that the surface-power received by the mike or your ears equals \(10^{-5}\) W/m² Solve : \(70= 10\lg \Large \frac{P_2}{10^{-12}}\)

OpenStudy (anonymous):

oh ok i understand that it's just that I don't know what P2 and P1 will be.

OpenStudy (anonymous):

what will they be?

OpenStudy (vincent-lyon.fr):

They will be the two powers you want to compare. P2 is ten times as big as P1 will become : P2 is P1 + 10dB P2 is a hundred times as big as P1 will become : P2 is P1 + 20dB P2 is a thousand times as big as P1 will become : P2 is P1 + 30dB etc.